# Surgical, Cheap, and Flexible: Mitigating False Refusal in Language Models via Single Vector Ablation

### v1

@@ 5.2 Fine-grained control with partial orthogonalization
The orthogonalization operation completely removes the component along the direction of the candidate true refusal vectors $\mathbf{v}_{i,l}$ from the candidate false refusal vectors $\mathbf{w}_{i,l}$. This operation maximally keeps the candidate true refusal vectors after the ablation, which makes the false refusal vector “softer” when removing the false refusal behaviour. Any feature in the pseudo-harmful queries associated with common features in the harmful queries will be regarded as a violation and will induce refusal, such as “how to cut off the heads of the fish” which is related to the violence feature, as we show in Fig 2. It is worth noting that there is no clear cut between what constitutes safe and unsafe queries, and no standard on how openly a model should answer. Therefore, the added flexibility of adapting the model to different refusal standards for different usage scenarios is important.
@@
Figure 2: Response of Llama2-7b-Chat to a XSTest-Safe sample under different $\lambda$ values for orthogonalization. The response openness increases as we lower the $\lambda$ value. The lower the $\lambda$ is, the less sensitive the model tends to answer the questions. The sensitivity level can be adjusted and calibrated by the user.
Through partial orthogonalization, we can control the distance between the candidate false refusal and true refusal vectors by controlling the coefficient $\lambda$, as introduced in Eq. 8. By decreasing the $\lambda$, the original false refusal vector is less modified by the orthogonalization, which leads to a stronger refusal removal when ablated. In Figure 2, as we decrease the $\lambda$ value the openness of the model response increases. When we set the $\lambda$ to $1$, the model refuses to give the instruction because it regards “cutting off the heads of fish” as not safe and not hygienic. As we lower the $\lambda$ to $0.8$, the model starts to give the instruction but cautions against it as it is harmful to the health of the fish and affects the flavor and texture. With a lower $\lambda$ value of $0.6$, the model is fully open to give the instructions without any resistance. Therefore, $\lambda$ enables us to control the safety sensitivity of the model in a very fine-grained way.
@@
To understand the impact of the $\lambda$ value comprehensively, we plot in Figure 3 the MMLU performance and the compliance rate on OR-Bench-Hard (OR CR), XSTest-Safe Hard (XSTest CR) and harmful data $D_{\text{harm}}$ (Harmful CR) across the four models we evaluated. As shown in Figure 3, the compliance rate on both harmful and harmless data increases as the $\lambda$ decreases. This shows the ability to granularly adjust the boundary between safe and unsafe concepts by adjusting $\lambda$. For Gemma-7b-It, Llama-3-8b-Ins and Llama2-7B-Chat model, the harmful query compliance rate increases faster than the model of Llama-2-13b-Chat. This means the false refusal vectors are closer to the true refusal vectors in these two models, which affects more the true refusal behaviour after the ablation. In comparison, the compliance rate on harmful data remains low when the $\lambda$ decreases for Llama2-13b-chat, while the compliance rate on pseudo-harmful data increases substantially. This indicates that false refusal and true refusal vectors we extracted from Llama2-13b-chat are well separated and disentangled in the beginning before the orthogonalization. Therefore, changing the $\lambda$ values has less impact on the true refusal.
The compliance rate curves give us an idea of how well the false refusal and true refusal are separated for the model and how to calibrate it. For models such as Gemma-7b-It, Llama2-7B-Chat and Llama-3-8b-Ins, staying in the range between $0.8$ and $1.0$ keeps the model relatively safe while lowering the $\lambda$ makes the model substantially more helpful, which proves a good trade-off between ‘harmlessness’ and ‘helpfulness’. As for Llama2-13-chat, it is safer to further lower the $\lambda$ to maximize the helpfulness of the model. One possible reason for such clear separation between false refusal and true refusal vectors is that Llama2-13-chat has undergone heavy safety finetuning which leads to the lowest compliance rate on pseudo-harmful data, reaching $5.5\%$ on ORB-H. This sets the boundary between a safe and unsafe query to an extreme level. Therefore, lowering the threshold for refusal will first relieve the pseudo-harmful ones before affecting the refusal to harmful queries which have stronger harmful-related features.
To understand the $\lambda$’s impact on general model capability, we also evaluate the model accuracy on MMLU in a 5-shot setting. As shown in Fig. 3, the MMLU accuracy is not affected by the $\lambda$ value across different models. This is achieved by filtering the candidate refusal vectors that have largely shifted the first token probability distribution before the ablation.
@@ 5.3 Refusal Vector Distribution
To better understand how the candidate true refusal and false refusal vectors are distributed in the transformer layers, we plot the refusal score changes across the token positions and layers when ablating the candidate refusal vectors in Figure 4. The y-axis is the relative token positions which are the positions of the post-instruction tokens such as [/INST]. The top row shows the refusal score changes when ablating the candidate true refusal vectors. From the second row, we plot the refusal score changes for the candidate false refusal vectors with increasing $\lambda$ value. Compared with the true refusal vector distribution, the false refusal vectors generally have a lower refusal score change when ablated. This is because the false refusal vectors are extracted from pseudo-harmful data, which is less likely to induce refusal compared to the true refusal vectors. We also see a position overlap between the two kinds of vectors, as shown in the first two rows in the figure. This indicates that the false and true refusal is activated at similar positions and simply removing one could affect the other one, which is shown in our earlier experiment in Table 1. As we increase the $\lambda$, the overall refusal score change is decreased, thus resulting in a weaker refusal removal effect. Furthermore, the distribution overlap between the candidate true refusal and false refusal vectors shrinks with increasing $\lambda$. For example, the most effective candidate true refusal vectors of Llama-3-8B-Ins are concentrated at layer $13$, while the most effective candidate false refusal vectors are mostly at layer $24$ as the $\lambda$ goes from 00 to $0.6$. Since we only select the most effective candidate refusal vector for the ablation, the final false refusal vector will be selected at different layer and token positions than the true refusal vector as we increase the $\lambda$, avoiding harming the true refusal behaviour.
@@
Figure 4: Refusal score changes when ablating the refusal vector extracted at certain layers and token positions. By increasing the value of $\lambda$, the refusal vectors have less impact on the model refusal behaviour.
@@
5.4 Does Adding True Refusal Vector Help?
@@
Arditi et al. (2024) show that the model can be made to refuse more by adding the refusal vector to the activation stream. It is potentially beneficial to apply both the false refusal vector ablation and true refusal vector addition to get a more useful model without harming the safety guard. To investigate the effectiveness of this approach, we evaluate the performance of Llama-2-7b-chat on harmful data, pseudo-harmful data and general tasks under each combination of the adding coefficient $\alpha$ and orthogonalization coefficient $\lambda$. Results are given in Figure 8. We refer readers to the Appendix C for results of Gemma-7b model. The baseline value below each plot shows the original model performance before the intervention. As shown in Figure 8, adding the true refusal vector can indeed make the model safer by inducing the refusal behaviour. In the region where $\alpha\geq 0.2$ and $\lambda\geq 0.6$ as highlighted in the figure, the model behaves relatively safe on the Jailbreakbench with a low compliance rate compared to the original model (baseline). This also leads to a low compliance rate on pseudo-harmful data such as OR-Bench-Hard. However, we see a degradation of general capacity measured by the accuracy on ARC-C, MMLU and perplexity on Wikitext. Therefore, the ablation operation is a more surgical approach than the addition, which was used by SCAN (An et al., 2024a), and indeed leads to a general performance degradation.
@@
However, the performance on TruthfulQA (Lin et al., 2022) has increased as the $\alpha$ increases. The BLUE score reaches $64.6$ under the setting of $\alpha=0.8,\lambda=0.8$, compared to the baseline value of $45.3$. As we inspect the model responses, adding the refusal vector tends to make the model more conservative or refusing to answer, which is favoured by the TruthfulQA questions. Based on above observation, it is important to pay attention to the general performance degradation when adding true refusal vector to make the model safer since it is not as surgical as the ablation operation.
@@ 7 Conclusion
In this work, we propose a surgical approach to mitigate the false refusal in language models via ablating a single vector from the activation stream of the transformer model. This is done by ablating a false refusal vector extracted from the mean model activation difference when prompting the model with pseudo-harmful and harmless data. To avoid compromising the safety guard against harmful queries, we apply orthogonalization to remove the projection of the false refusal vector onto the true refusal vector. We also show that fine-grained calibration of the safety guard can be achieved by partial orthogonalization and adjusting the strength of the projection removal. This enables more flexibility for adjusting the response openness to ambiguous and sensitive queries that are up to the user’s judgment. The surgical characteristic of the ablation operation keeps the modified model’s original general capabilities and requires no additional inference cost via direct model weight modification.
@@
We show that the false refusal vector can be extracted using a small sample of pseudo-harmful and harmless queries. However, the quality of the vector is not optimal as we mainly focused on using pseudo-harmful samples from OR-Bench-Hard (Cui et al., 2024b), which showed good performance at mitigating false refusal. A better data curation strategy and sampling technique could further improve the effectiveness of the vector, which we leave for future work. To select the refusal vector, we adopt the refusal score metrics from previous work (Arditi et al., 2024) which looks at the refusal-related tokens at the first token position. Such evaluation metrics can be further improved, such as considering sequence probabilities, for better refusal vector selection.
@@ Appendix A System Prompt
Appendix B Evaluation on Safety and Refusal Output
B.1 String Matching
@@
B.2 Human and LLM Evaluation
Although the models tend to refuse with these keywords, we manually checked part of the result. We found out that while string matching could detect most of the refusal cases on pseudo-harmful datasets, there is a minor gap between string matching and human annotation results on harmful datasets. As such, we utilized Llama-Guard-3-8B (Llama Team, 2024) for safety evaluation in our main experiment in Table 2, 3 and further checked the result with two human annotators. We use string matching for other experiments for efficiency and refusal-changing trend analysis. For general refusal trend analysis, string matching shows similar results as LlamaGaurd in our preliminary experiments.
Appendix C Refusal Vector Addition
@@
Appendix D Fine-grained Control on Response Openness
@@
This section provides further details on how varying the orthogonalization parameter $\lambda$ allows fine-grained control over the openness of model responses to potentially sensitive or harmful prompts, with examples in Fig 9 to 12. As $\lambda$ decreases from 1 (full orthogonalization) to 0 (no orthogonalization), the responses tend to become more direct in addressing the prompt’s request.

### v2

@@ 5.2 Fine-grained control with partial orthogonalization
The orthogonalization operation removes the component along the direction of the candidate true refusal vectors $\mathbf{v}_{i,l}$ from the candidate false refusal vectors $\mathbf{w}_{i,l}$. This operation keeps the candidate true refusal vectors after the ablation, which makes the false refusal vector “softer” when removing the false refusal behavior. Any feature in the pseudo-harmful queries associated with common features in the harmful queries will be regarded as a violation and will induce refusal, such as “how to cut off the heads of the fish” which is related to the violence feature, as we show in Figure 2.
@@
Figure 2: Response of Llama2-7b-Chat to a XSTest-Safe samples under different $\lambda$ values. The response openness increases as we lower the $\lambda$ value. The lower the $\lambda$ is, the less sensitive the model tends to answer the questions. The sensitivity level can be adjusted by the user.
s It is worth noting that there is no clear cut between what constitutes safe and unsafe queries, and no standard on how compliant the answer should be. Therefore, the added flexibility of adapting the model to different refusal standards for different usage scenarios is important.
Through partial orthogonalization, we can control the distance between the candidate false refusal and true refusal vectors by controlling the coefficient $\lambda$, as introduced in Eq. 8. By decreasing $\lambda$, the original false refusal vector is less modified by the orthogonalization, which leads to a stronger refusal removal when ablated. In Figure 2, as we decrease $\lambda$, model responses become more compliant. When we set $\lambda$ to $1$, the model refuses to give the instruction because it regards “cutting off the heads of fish” as not safe and not hygienic. As we lower $\lambda$ to $0.8$, the model starts to give a compliant answer but cautions against it as it is harmful to the health of the fish and affects the flavor and texture. With a lower $\lambda$ value of $0.6$, the model is fully compliant. Therefore, $\lambda$ enables us to control the safety sensitivity of the model in a very fine-grained way.
@@
To understand the impact of the $\lambda$ value comprehensively, we plot in Figure 3 the MMLU performance and the compliance rate on OR-Bench-Hard (OR CR), XSTest-Safe Hard (XSTest CR) and harmful data $D_{\text{harm}}$ (Harmful CR) across the four models we evaluated. As shown in Figure 3, the compliance rate on both harmful and harmless data increases as the $\lambda$ decreases. This shows the ability to granularly adjust the boundary between safe and unsafe concepts by adjusting $\lambda$. For Gemma-7b-It, Llama-3-8b-Ins and Llama2-7B-Chat model, the harmful query compliance rate increases faster than the model of Llama-2-13b-Chat. This means the false refusal vectors are closer to the true refusal vectors in these two models, which affects more the true refusal behavior after the ablation. In comparison, the compliance rate on harmful data remains low when the $\lambda$ decreases for Llama2-13b-chat, while the compliance rate on pseudo-harmful data increases substantially. This indicates that false refusal and true refusal vectors we extracted from Llama2-13b-chat are well separated and disentangled in the beginning before the orthogonalization. Therefore, changing the $\lambda$ values has less impact on the true refusal.
The compliance rate curves give us an idea of how well the false refusal and true refusal are separated for the model and how to calibrate it. For models such as Gemma-7b-It, Llama2-7B-Chat and Llama-3-8b-Ins, staying in the range between $0.8$ and $1.0$ keeps the model relatively safe on harmful queries. Setting $\lambda$ as $0.8$ makes the model substantially more helpful, which proves a good trade-off between ‘harmlessness’ and ‘helpfulness’. As for Llama2-13B-chat, it is safer to further lower the $\lambda$ to maximize the helpfulness of the model. One possible reason for such clear separation between false refusal and true refusal vectors is that Llama2-13B-chat has undergone heavy safety finetuning which leads to the lowest compliance rate on pseudo-harmful data, reaching $5.5\%$ on ORB-H. This sets the boundary between a safe and unsafe query to an extreme level. Therefore, lowering the threshold for refusal will first affect the pseudo-harmful ones before affecting refusal to harmful queries, which have stronger harmful-related features.
To understand $\lambda$’s impact on general model capability, we also evaluate the model accuracy on MMLU in a 5-shot setting. As shown in Figure 3, the MMLU accuracy is not affected by changing $\lambda$ across different models. The fact that performance is not affected is because we filtered the candidate refusal vectors that have largely shifted the first token probability distribution before the ablation.
@@ 5.3 Refusal Vector Distribution
To better understand how the candidate true refusal and false refusal vectors are distributed in the transformer layers, we plot the refusal score changes across the token positions and layers when ablating the candidate refusal vectors in Figure 4. The y-axis is the relative token positions which are the positions of the post-instruction tokens such as [/INST]. The top row shows the refusal score changes when ablating the candidate true refusal vectors. From the second row, we plot the refusal score changes for the candidate false refusal vectors with increasing $\lambda$ value. Compared with the true refusal vector distribution, the false refusal vectors generally have a lower refusal score change when ablated. This is because the false refusal vectors are extracted from pseudo-harmful data, which is less likely to induce refusal compared to the true refusal vectors. We also see a position overlap between the two kinds of vectors, as shown in the first two rows in the figure. This indicates that the false and true refusal is activated at similar positions and simply removing one could affect the other one, which is shown in our earlier experiment in Table 1. As we increase $\lambda$, the overall refusal score changes less, thus resulting in a weaker refusal removal effect. Furthermore, the distribution overlap between the candidate true refusal and false refusal vectors shrinks with increasing $\lambda$. For example, the most effective candidate true refusal vectors of Llama-3-8B-Ins are concentrated at layer $13$, while the most effective candidate false refusal vectors are mostly at layer $24$ as $\lambda$ increases from 00 to $0.6$. Since we only select the most effective candidate refusal vector for the ablation, the final false refusal vector will be selected at different layer and token positions than the true refusal vector as we increase the $\lambda$, avoiding harming the true refusal behavior.
@@
Figure 4: Refusal score changes when ablating the true (row 1) and false (row 2,3) refusal vectors extracted at certain layers and token positions. By increasing the value of $\lambda$, the refusal vectors have less impact on the model refusal behaviour.
@@
5.4 Does Adding True Refusal Vector Help Keeping Model Safe?
@@
Arditi et al. (2024) show that the model can be made to refuse more by adding the refusal vector to the activation stream. It is potentially beneficial to apply both the false refusal vector ablation and true refusal vector addition to get a more useful model without harming the safety guard. To investigate the effectiveness of this approach, we evaluate the performance of Llama-2-7b-chat on harmful data, pseudo-harmful data and general tasks under each combination of the adding coefficient $\alpha$ and orthogonalization coefficient $\lambda$. Results are given in Figure 9. We refer readers to Appendix E for results of Gemma-7b model. The baseline value below each plot shows the original model performance before the intervention. As shown in Figure 9, adding the true refusal vector can indeed make the model safer by inducing more refusal behavior. In the region where $\alpha\geq 0.2$ and $\lambda\geq 0.6$ as highlighted in the figure, the model behaves relatively safe on Jailbreakbench with a low compliance rate compared to the original model (baseline). This also leads to a low compliance rate on pseudo-harmful data such as OR-Bench-Hard. However, we see a degradation of general capacity measured by the accuracy on ARC-C, MMLU and perplexity on Wikitext. Therefore, the ablation operation is a more surgical approach than the addition, which was used by SCAN (An et al., 2024a) leading to general performance degradation, while our method does not.
@@
However, the performance on TruthfulQA (Lin et al., 2022) increases as $\alpha$ increases. The BLUE score reaches $64.6$ under the setting of $\alpha=0.8,\lambda=0.8$, compared to the baseline value of $45.3$. As we inspect the model responses, adding the refusal vector tends to make the model more conservative and more likely to refuse to answer, which is favoured by the TruthfulQA questions. Based on the above observation, it is important to pay attention to the general performance degradation when adding true refusal vector to make the model safer, since it is not as surgical as the ablation operation. Therefore, we propose ablation as a mitigation approach versus the addition in Shi et al. (2024).
@@ 7 Conclusion
In this work, we proposed a surgical approach to mitigate the false refusal in language models via ablating a single vector from the activation stream of the transformer model. We ablated a false refusal vector extracted from the mean model activation difference when prompting the model with pseudo-harmful and harmless data. To avoid compromising model safety on harmful queries, we applied orthogonalization to remove the projection of the false refusal vector onto the true refusal vector. We also showed that fine-grained calibration of the safety guard can be achieved by partial orthogonalization and adjusting the strength of the projection removal. This enables more flexibility for adjusting the response compliance level to ambiguous and sensitive queries that are up to the user’s judgment. The surgical characteristic of the ablation operation keeps the modified model’s original general capabilities and requires no additional inference cost via direct model weight modification.
@@
We show that the false refusal vector can be extracted using a small sample of pseudo-harmful and harmless queries. However, the quality of the vector can be further improved by constructing more diverse pseudo-harmful samples, as we mainly focused on using pseudo-harmful samples from OR-Bench-Hard (Cui et al., 2024b). A better data curation strategy and sampling technique could further improve the effectiveness of the vector, which we leave for future work. To select the refusal vector, we adopt the refusal score metrics from previous work (Arditi et al., 2024) which looks at the refusal-related tokens at the first token position. Such evaluation metrics can be further improved, such as considering sequence probabilities, for better refusal vector selection.
@@ Appendix A System Prompt
Appendix B Workflow
To give a clear view of how our vector ablation method works, we give a visualisation of the general workflow in Figure 7. We first collect (pseudo) harmful and harmless query pairs to extract (false) refusal vector using the diff-in-means method. Then, we apply orthogonalization between the true and false refusal vectors, to minimize the impact on the true vector when ablating the false refusal vector. Finally, we ablate the orthogonalized false refusal vector from the model, either during inference time or directly on the model weights. Section D provides more details about the equivalence between the two.
Figure 7: Our refusal vector extraction and ablation pipeline. It applies to both true and false refusal vector extraction and ablation.
Appendix C Evaluation on Safety and Refusal Output
C.1 String Matching
@@
C.2 Human and LLM Evaluation
Although the models tend to refuse with these keywords, we manually checked part of the result. We found out that while string matching could detect most of the refusal cases on pseudo-harmful datasets, there is a minor gap between string matching and human annotation results on harmful datasets. As such, we utilized Wildguard (Han et al., 2024) for safety evaluation in our main experiment in Table 2, 3. We further checked the result in Table 3 with two human annotators for a fair comparison with the baseline. We use string matching for other experiments for efficiency and refusal-changing trend analysis. For general refusal trend analysis, string matching shows similar results as WildGuard in our preliminary experiments.
Appendix D Equivalence to model editing through weight orthogonalization
The vector ablation can be applied during the inference time (Activation Steering) or before the inference time (Model Editing), as the vector we extracted is fixed and independent of the input. We directly apply the vector ablation on the weight matrix, mitigating the false refusal problem on the edited model, which requires no intervention during its inference time. As the new weight matrix through linear transformation keeps the original matrix size, the memory and inference time are kept. We give a detailed explanation below:
In the Transformer model, each Attention (Att) and Feedforward Neural Network (FNN) block writes its output to the residual stream. To prevent the Att and FNN blocks from representing the vector $\mathbf{r}$ we want to ablate, we can apply the following transformation during inference time:
$\mathbf{X}_{output}\rightarrow\mathbf{X}_{output}-\mathbf{r}\mathbf{r}^{T}\mathbf{X}_{output}$ (9)
where $\mathbf{X}_{output}$ is the output from either the Att or FNN block.
Given that the representation in each Att/FNN layer passes through a weight matrix $\mathbf{W}$ before being written to the residual stream: $\mathbf{X}_{output}=\mathbf{W}\mathbf{X}_{pre}$, ablating on $\mathbf{X}_{output}$ is equivalent to directly ablating the weight matrix $\mathbf{W}$:
$\mathbf{X}_{output}-\mathbf{r}\mathbf{r}^{T}\mathbf{X}_{output}$ $=\mathbf{W}\mathbf{X}_{pre}-\mathbf{r}\mathbf{r}^{T}(\mathbf{W}\mathbf{X}_{pre})$
$=(\mathbf{W}-\mathbf{r}\mathbf{r}^{T}\mathbf{W})\mathbf{X}_{pre}$
$=\mathbf{W}^{\prime}\mathbf{X}_{pre}$ (10)
, where $\mathbf{X}_{pre}$ is the representation before the final linear layer in the Att/FNN block.
The new matrix $\mathbf{W}^{\prime}$ has the same dimensions as the original matrix $\mathbf{W}$, resulting in no additional memory or inference cost. We also refer readers to Appendix E in Arditi et al. (2024), where the authors also provided detailed proof of this equivalence.
Appendix E Refusal Vector Addition
@@
Appendix F Comparison to more baseline methods
Here we compare our method to other previous works on safe-guarding or false refusal mitigation, including training-based methods such as DRO (Zheng et al., 2024), and training-free methods such as Self-CD (Shi et al., 2024), SafeDecoding(Xu et al., 2024) and SCANS (Cao et al., 2024). All baseline performance numbers are directly taken from Cao et al. (2024). As we discussed in the main text, SCANS achieves a higher compliance rate on XSTest-Safe with a cost of a higher compliance rate on XSTest-Unsafe and lower performance on MMLU and Wikitext perplexity. Among other methods, the training-based method DRO achieves the lowest compliance rate on XSTest-Safe, indicating a serious false refusal problem. For the training-free approaches, both SafeDecoding and Self-CD achieve a higher CR on the Unsafe part of XSTest and a lower/equal CR on the Safe part of XSTest.
XSTest-U $\downarrow$ XSTest-S $\uparrow$ SafeDecoding 0.5 75.2 Self-CD 2.5 85.2 DRO 1.5 58.5 SCAN 6.5 91.8 Ours 0 85.2
Table 4: Compliance rate of Llama2-7B-Chat after intervention on XSTest-Unsafe and XSTest-Safe.
Appendix G Fine-grained Control on Response Openness
@@
This section provides further details on how varying the orthogonalization parameter $\lambda$ allows fine-grained control over the openness of model responses to potentially sensitive or harmful prompts, with examples in Fig 10 to 13. As $\lambda$ decreases from 1 (full orthogonalization) to 0 (no orthogonalization), the responses tend to become more direct in addressing the prompt’s request.